### v1

# BAMBINO-LM: (Bilingual-)Human-Inspired Continual Pretraining of BabyLM

Abstract
Children from bilingual backgrounds benefit from interactions with parents and teachers to re-acquire their heritage language. In this paper, we investigate how this insight from behavioral study can be incorporated into the learning of small-scale language models. We introduce BAMBINO-LM, a continual pretraining strategy for BabyLM that uses a novel combination of alternation and PPO-based perplexity reward induced from a parent Italian model. Upon evaluation on zero-shot classification tasks for English and Italian, BAMBINO-LM improves the Italian language capability of a BabyLM baseline. Our ablation analysis demonstrates that employing both the alternation strategy and PPO-based modeling is key to this effectiveness gain. We also show that, as a side effect, the proposed method leads to similar degradation in L1 effectiveness as human children would have had in an equivalent learning scenario.
BAMBINO-LM: (Bilingual-)Human-Inspired Continual Pretraining of BabyLM
Zhewen Shen1 Aditya Joshi 1 Ruey-Cheng Chen2 1 University of New South Wales, Sydney, Australia 2 Canva, Sydney, Australia zhewen.shen@student.unsw.edu.au, aditya.joshi@unsw.edu.au, rcchen@canva.com
1 Introduction
The recently held BabyLM challenge (Warstadt et al., 2023) explores pretraining of language models using a constrained dataset analogous to the linguistic exposure of a 13-year-old English-speaking child. In this paper, we extend the BabyLM challenge to a bilingual setting, drawing inspiration from parent-child interactions in heritage language acquisition Lohndal et al. (2019). Immigrant children in western societies, who may have acquired their home language at a young age, can sometimes need to re-acquire the same language during the school years when the language becomes a minority. These heritage speakers typically benefit from an extended exposure to the minority language at home or in the community, owing largely to feedback and stimuli provided by parents and family members Montrul (2010). This observation about child bilingualism is in line with the behaviorist theory for child language development Demirezen (1988). Inspired by this line of work, we ask the following research question:
Can a small-scale language model trained on the majority language (e.g. English) be continually pretrained on the minority language, leveraging the feedback of a second model that is fluent in the latter language?
To address this question, we introduce ‘Bilingual language Acquisition Modeling Based on INterleaved Optimization of Language Models (BAMBINO-LM)’, a novel continual pretraining strategy that uses a combination of alternation and proximal policy optimization (PPO) using a reward from a second model playing the parent role (i.e., a large language model pretrained in the minority language). We experiment with BabyLM trained on English, and continually pretrain this model on an assumed second language, Italian. In its connection to cognitive processing, our work makes the following contributions:
  • BAMBINO-LM draws inspiration from bilingual language acquisition and learns from interactions with a second model by incorporating a perplexity-based reward for language model pretraining.
  • We show that BAMBINO-LM can acquire Italian to a reasonable degree with some expected degradation in its English capability. The findings hints at a common learning trajectories for second language acquisition shared by language models and humans.
2 Related Work
Pretraining small-scale language models is an emerging field that has garnered some interest from the language acquisition community. BabyBERTa Huebner et al. (2021) is an early adaptation to this scenario. Warstadt et al. (2023) introduce the BabyLM challenge to provide an atypically small dataset for benchmarking small-scale language models. This shared task enables research in not only language acquisition but sample-efficient pretraining.
Our work is primarily focused on continual pretraining and conducted in a setup similar to Yadavalli et al. (2023), where a tiered first/second language acquisition process is attempted. Samuel (2023) also experiments with a teacher-student setting but only tests the approach on English tasks. Evanson et al. (2023) is another closely related work, which investigates the learning trajectory of large-scale language models by probing their syntactic and semantic capabilities at each step. Contrasting GPT-2 models with children, the study shows that the linguistic skills are acquired sequentially in a similar order.
Reinforcement learning from human feedback (RLHF) utilizes human preferences for serving reward signals when dealing with sparse training labels Christiano et al. (2017); Stiennon et al. (2020), and has been shown successful for generative tasks such as dialogues and summarization. This approach is further extended in Bai et al. (2022) by using AI feedback (RLAIF) to remove the dependency on human preference data, leading to better scalability and signal availability. Our work departs from these prior efforts in the way a perplexity reward is induced and blended into the learning process.
3 Methods
Figure 1: Architecture of BAMBINO-LM.
Figure 1 shows the two phases of BAMBINO-LM. The learning phase involves continual pretraining a small-scale language model (baby model $\mathcal{B}$) whose initial pretraining was originally done on English data, while the feedback phase involves interactions with the Italian language model (parent model $\mathcal{P}$). During the learning phase, pretraining for $\mathcal{B}$ is continued by employing causal language modeling on Italian data. Causal language modeling (CLM), also known as next token prediction, is a standard technique to train a decoder-only model. The objective is defined as follows:
$\mathcal{L}_{\operatorname{CLM}}=-\frac{1}{|x|}\sum_{i=0}^{|x|}\log\mathbb{P}(x_{t}\mid x_{0},\ldots,x_{t-1}).$
There are two architectural innovations in BAMBINO-LM:
Feedback phase based on PPO
We construct prompt $x$ by selecting the first $k$ tokens from the training example, and solicit output $y_{\mathcal{B}}=\mathcal{B}(x)$ from the baby model. We then use Proximal Policy Optimization (PPO) where $\mathcal{B}$’s parameters are updated according to a clipped surrogate objective Schulman et al. (2017). This objective moderates updates to the policy, facilitating stable and efficient learning by incorporating a clipping mechanism. Its definition is given as follows:
$\mathcal{L}_{\text{PPO}}=\mathbb{E}_{t}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\operatorname{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right].$
Note that the advantage function $\hat{A}_{t}$ is of the form:
$\hat{A}_{t}=R+\gamma V(s_{t+1})-V(s_{t}).$
The reward $R$ for the advantage function is then calculated using the following function:
$R(y_{\mathcal{B}})=\frac{\alpha}{\beta(\operatorname{PPL}_{\mathcal{P}}(y_{\mathcal{B}})-\tau)},$ (1)
where $\alpha$ and $\beta$ are parameters, $\operatorname{PPL}_{\mathcal{P}}$ represents the perplexity of the parent model $\mathcal{P}$ for the sequence $y_{\mathcal{B}}$, and $\tau$ is a threshold value for perplexity. We use the following formulation of perplexity:
$\operatorname{PPL}(x)=\exp\left[\sum_{i=0}^{|x|}\log\mathbb{P}(x_{t}\mid x_{0},\ldots,x_{t-1})\right].$
Alternating run
We adopt an alternating run strategy between the learning and feedback phases, which is summarized in Algorithm 1. The rationale behind this is two-fold: 1) this strategy simulates frequent interactions between a child and its parent through dialogues, which has been our main motivation behind this study; 2) using multiple rewards is shown beneficial for reinforcement learning Dann et al. (2023). To expand on the second point, our findings further suggest that using perplexity as a reward can lead to exploitation when baby model $\mathcal{B}$ attempts to produce similar utterances to those coming from parent model $\mathcal{P}$. Without this strategic alternation between CLM and PPO, the pretraining tends to produce undesirable behaviours such as repeating words.
Algorithm 1 BAMBINO-LM Training.
1:procedure Train($\mathcal{D},\mathcal{B},\mathcal{P}$)
2:     input: pretraining dataset $\mathcal{D}$, baby model $\mathcal{B}$, and parent model $\mathcal{P}$.
3:     $r_{\operatorname{CLM}},r_{\operatorname{PPO}}\leftarrow 10,2$
4:     $r\leftarrow r_{\operatorname{CLM}}+r_{\operatorname{PPO}}$
5:     for $i,x\in\operatorname{enumerate}(\mathcal{D}$do
6:         if $i\%r<r_{\operatorname{CLM}}$ then
7:              perform CLM step
8:         else
9:              $y_{\mathcal{B}}\leftarrow\mathcal{B}(x[1..k])$
10:              reward $\leftarrow R(y_{\mathcal{B}})$
11:              perform PPO step
12:         end if
13:     end for
14:end procedure
4 Experiment Setup
Mimicking their process to create the BabyLM challenge corpus (Warstadt et al., 2023), we created an Italian dataset that is comparable in size to the strict-small track of the challenge, and perform identical preprocessing [^1] https://github.com/babylm/babylm_data_preprocessing; Accessed on 13th May, 2024.. Table 1 shows the statistics of the Italian language dataset.
For the choice of the baby model, we use English baseline OPT-125m Zhang et al. (2022) model for the strict-small track provided by the BabyLM organizers. For the parent model, we use gpt2-small-italian model by de Vries and Nissim (2021). Using the Italian dataset described above, we conduct continual pretraining over 10 epochs, consisting of 10 learning phase steps followed by 2 feedback phase steps. We use $k=5$ to solicit the first few tokens for prompting the baby model. All models are trained using HuggingFace’s transformer (Wolf et al., 2020) and trl (von Werra et al., 2020) library.
Dataset %
CHILDES (MacWhinney, 2000) 2.23
DailyDialog (Li et al., 2017) 4.45
QED (Abdelali et al., 2014) 11.86
OpenSubtitles Lison and Tiedemann (2016) 27.58
Standardised Project Gutenberg Corpus Gerlach and Font-Clos (2020) 16.19
Children’s Story  [^2] https://www.gutenberg.org/ebooks/bookshelf/353 18.57
Wikipedia  [^3] https://dumps.wikimedia.org/itwiki/ 19.10
Table 1: Italian dataset used for continual pretraining.
For downstream tasks, we use four Italian language tasks in UINAUIL Basile et al. (2023) and four English language tasks in GLUE Wang et al. (2018). The tasks were selected primarily based on computational constraints for the project. We also include BLiMP Warstadt et al. (2020) for that it is used in the original BabyLM challenge. All tasks are conducted in a zero-shot classification setting.
5 Results
Task / Model BabyLM BAMBINO-LM
UINAUIL
HaSpeeDe 0.4774 0.4592
IronITA 0.4966 0.5516
SENTIPOLC 0.1575 0.4050
Textual Entailment 0.4950 0.5525
Average 0.3416 0.4613
GLUE
MNLI 0.3472 0.3530
MNLI-MM 0.3483 0.3521
RTE 0.5271 0.5199
SST2 0.5034 0.5241
Average 0.4315 0.4373
BLiMP
Average 0.6255 0.5503
Table 2: Comparison of BAMBINO-LM with BabyLM.
Table 2 shows a significant improvement in Italian downstream tasks for BAMBINO-LM as compared with the BabyLM baseline. Specifically, we achieve an average improvement of 0.1197 $(0.3416\to 0.4613)$ without substantial differences in English classification tasks. However, we notice an expected decrease of 0.0752 $(0.6255\to 0.5503)$ in the English language BLiMP dataset. These observations are in line with Yadavalli et al. (2023) which show that native child-directed speech can lead to negative cross-lingual transfer and impede L2 acquisition depending on the choice of L1.
In Table 3 we examine two ablated versions of our model: (a) w/o PPO: Trained solely on the CLM objective with no feedback phase; (b) w/o alternating: Use BAMBINO-LM with no alternating runs. Instead, it trains with the CLM objective for the first $85\%$ of each epoch and then switches to the PPO objective for the remaining $15\%$.
Removing the interactive feedback mechanism (w/o PPO) and the alternating strategy (w/o alternating) significantly decreases Italian performance compared to our primary model. On UINAUIL tasks, the average score drops from 0.4613 to 0.4000 (w/o PPO) and 0.3513 (w/o alternating). However, we do not observe significant improvements in performance in both the English language task sets (GLUE and BLiMP). For GLUE tasks, the average scores remain consistent, with 0.4373 for BAMBINO-LM, 0.4375 for w/o PPO, and 0.4357 for w/o alternating. On the BLiMP dataset, the average scores are 0.5503 for BAMBINO-LM and 0.5554 for w/o PPO.
Task / Model BAMBINO-LM w/o PPO BAMBINO-LM w/o alternating
UINAUIL
HaSpeeDe 0.4798 0.4925
IronITA 0.4966 0.4989
SENTIPOLC 0.2775 0.1580
Textual Entailment 0.5500 0.5500
Average 0.4000 0.3513
GLUE
MNLI 0.3540 0.3522
MNLI-MM 0.3502 0.3545
RTE 0.5343 0.5271
SST2 0.5115 0.5092
Average 0.4375 0.4357
BLiMP
Average 0.5554 0.5268
Table 3: Results of the ablation experiments.
These results indicate that PPO modeling and alternating runs are both crucial for improving the bilingual ability of BAMBINO-LM without negatively impacting English performance. Furthermore, the lack of significant changes in English scores reinforces that these strategies enhance bilingual capabilities without compromising performance on existing benchmarks.
6 Conclusion
This paper introduces BAMBINO-LM, a continual pretraining strategy mimicking the process of second language acquisition under an interactive setting. BAMBINO-LM uses a two-phase approach: it incorporates reward from a parent Italian model into a PPO-based mechanism, and alternates this procedure together with causal language modeling on top of Italian text. Our experiments demonstrate systematic improvement in Italian with a marginal but expected decrease in English, which echoes the past results in second language acquisition for large language models Evanson et al. (2023). These findings highlight the efficacy of our approach in enhancing bilingual capabilities while maintaining performance in the original language.
In future work, we aim to explore the effect of alternative metrics and different reward learning mechanisms that better align with human feedback behaviors. This also includes exploring rewards that capture linguistic quality and provide direct, “constructive” corrections to the model output which is commonly known as an effective learning strategy for language development.
Limitations
The approach relies on availability of a base model in a language, English in our case. Although we download Italian language datasets from known Italian sources, we do not explicitly validate the language of the text. We use the PPO model as is, and do not experimentally tune its parameters. Similarly, using perplexity as a metric for computing rewards may not be the optimal solution, as perplexity itself is influenced by many factors of the parent model.
Ethics Statement
The paper uses publicly available datasets for training and evaluation that do not possess known harms. The evaluative tasks are typical language learning tasks. However, the resultant models are not tested for harmful or biased content.